Write an equation for a line passing through the points (t, 2b) and (c, 3b)?
Determine whether the graph of the given equation are parallel,perpendicular, or neither y=x+11
y=-x+2. Please help me. I'm confused. Thank you so much

Answer by rothauserc(4718)

The standard form of a line is y = mx +e where m is the slope and e is the y intercept
we use the points (t, 2b) and (c, 3b) to calculate the slope
m = (3b -2b) / (c -t) = b / (c-t)
therefore y = bx/(c-t) +e
note that the slope of y = x+11 is 1 and
the slope of y = -x+2 is -1
two lines are perpendicular if the product of their slopes are -1
these two lines are perpendicular